# Exact scattering cross section for lattice-defect scattering of phonons

### v1

@@ IV Summary and conclusions
In summary, we have shown rigorously how the mode-dependent scattering cross sections and rates in a lattice can be determined using the extended AGF method. (Ong, 2018a, b) This AGF-based approach allows us to treat scattering problems in which additional degrees of freedom from an external scatterer are present. Its utility is demonstrated with two examples: (1) the monoatomic harmonic chain with a single isotopic impurity, and (2) the (10,10) CNT with an encapsulated C60 molecule. In the first example, we derive the exact analytical expressions for the scattering cross section and rate. To first order in pertubation, the latter is identical to the result in Ref. (Allen and Nghiem, 2022). In the second example, we illustrate how the scattering cross section varies with the properties of the phonon mode (phonon branch and frequency), and show how phonon modes with radial atomic motion are more strongly scattered by the C60 molecule.
@@ Appendix A The extended Atomistic Green’s Function method
where $\boldsymbol{H}_{\text{C}}$, and $\boldsymbol{H}_{\text{CL}}$ ($\boldsymbol{H}_{\text{CR}}$) are respectively the mass-normalized force-constant submatrices corresponding to the device region, and the coupling between the device region and the semi-infinite left (right) lead. We can associate each layer in Fig. 1 with a block row in $\mathbf{H}$. The block row submatrices $\boldsymbol{H}_{\alpha}^{00}$ and $\boldsymbol{H}_{\alpha}^{01}$, where $\alpha=\text{L}$ and $\alpha=\text{R}$ for the left and right lead, respectively, describe the lead phonons. If we set the layers to be large enough so that only adjacent layers can couple, then $\boldsymbol{H}_{\alpha}^{00}$ denotes to the intralayer force-constant matrix for each layer while $\boldsymbol{H}_{\alpha}^{01}$ ($\boldsymbol{H}_{\alpha}^{10}$) denotes the interlayer harmonic coupling between each layer and the layer to its right (left) in the lead. Here, $\alpha$ is used as the dummy variable for distinguishing the leads, with $\alpha=\text{L}$ and $\alpha=\text{R}$ corresponding to the left and right lead, respectively. We also maintain the conceptual distinction between the left-lead and right-lead force-constant matrices of $\boldsymbol{H}_{\alpha}^{00}$ and $\boldsymbol{H}_{\alpha}^{01}$ even though $\boldsymbol{H}_{\text{L}}^{00}=\boldsymbol{H}_{\text{R}}^{00}$ and $\boldsymbol{H}_{\text{L}}^{01}=\boldsymbol{H}_{\text{R}}^{01}$ in my method for computing the scattering cross section.
@@
In principle, the system dynamics are determined by the infinitely large $\mathbf{H}$ in Eq. (15). However, for the effective dynamics at a fixed frequency $\omega$, the lattice dynamics problem becomes more tractable as we need only to project the dynamics onto a finite portion of the system, (Wang et al., 2008; Mingo, 2009) corresponding to layers $n=0$ to $N$ in Fig. 1, to determine phonon scattering by the device region. Hence, we can use the submatrices in Eq. (15) to construct the effective harmonic matrix for this subsystem (Wang et al., 2008)
@@
where $\boldsymbol{H}_{\text{L}}^{\prime}=\boldsymbol{H}_{\text{L}}^{00}+\boldsymbol{H}_{\text{L}}^{10}\boldsymbol{g}_{\text{L},-}^{\text{ret}}\boldsymbol{H}_{\text{L}}^{01}$ and $\boldsymbol{H}_{\text{R}}^{\prime}=\boldsymbol{H}_{\text{R}}^{00}+\boldsymbol{H}_{\text{R}}^{01}\boldsymbol{g}_{\text{R},+}^{\text{ret}}\boldsymbol{H}_{\text{R}}^{10}$ correspond to the left and right edge, respectively while $\boldsymbol{H}_{\text{C}}^{\prime}=\boldsymbol{H}_{\text{C}}$ and $\boldsymbol{H}_{\text{CL/CR}}^{\prime}=\boldsymbol{H}_{\text{CL/CR}}=(\boldsymbol{H}_{\text{LC/RC}}^{\prime})^{\dagger}$. The retarded surface Green’s functions $\boldsymbol{g}_{\text{L},-}^{\text{ret}}(\omega)$ and $\boldsymbol{g}_{\text{R},+}^{\text{ret}}(\omega)$ are given by
$\boldsymbol{g}_{\alpha,-}^{\text{ret}}=[(\omega^{2}+i0^{+})\boldsymbol{I}_{\alpha}-\boldsymbol{H}_{\alpha}^{00}-\boldsymbol{H}_{\alpha}^{10}\boldsymbol{g}_{\alpha,-}^{\text{ret}}\boldsymbol{H}_{\alpha}^{01}]^{-1}$ (17a)
$\boldsymbol{g}_{\alpha,+}^{\text{ret}}=[(\omega^{2}+i0^{+})\boldsymbol{I}_{\alpha}-\boldsymbol{H}_{\alpha}^{00}-\boldsymbol{H}_{\alpha}^{01}\boldsymbol{g}_{\alpha,+}^{\text{ret}}\boldsymbol{H}_{\alpha}^{10}]^{-1}$ (17b)
Physically, Eq. (17a) denotes the retarded surface Green’s function for a decoupled semi-infinite lattice extending infinitely to the left (denoted by the ‘-’ in the subscript of $\boldsymbol{g}_{\alpha,-}^{\text{ret}}$) while Eq. (17b) denotes the corresponding surface Green’s function for a decoupled semi-infinite lattice extending infinitely to the right (denoted by the ‘+’ in the subscript of $\boldsymbol{g}_{\alpha,+}^{\text{ret}}$). In addition, the advanced surface Green’s functions can be obtained from the Hermitian conjugates of Eq. (17b), i.e., $\boldsymbol{g}_{\alpha,-}^{\text{adv}}=(\boldsymbol{g}_{\alpha,-}^{\text{ret}}){}^{\dagger}$ and $\boldsymbol{g}_{\alpha,+}^{\text{adv}}=(\boldsymbol{g}_{\alpha,+}^{\text{ret}})^{\dagger}$.
@@
The next step is the computation of the advanced and retarded Bloch matrices (Ando, 1991; Khomyakov et al., 2005; Ong and Zhang, 2015) of the left and right lead, $\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(+)$ and $\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(-)$, which describe the bulk translational symmetry of the layers along the direction of the heat flux and can be computed directly from the formulas:
$\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(+)=\boldsymbol{g}_{\alpha,+}^{\text{adv/ret}}\boldsymbol{H}_{\alpha}^{10}$ (19a)
$\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(-)^{-1}=\boldsymbol{g}_{\alpha,-}^{\text{adv/ret}}\boldsymbol{H}_{\alpha}^{01}$ (19b)
The bulk eigenmodes for the lead, which describe the mode-dependent atomic displacement in each layer at frequency $\omega$, can be determined directly from the Bloch matrices:
$\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(+)\boldsymbol{U}_{\alpha}^{\text{adv/ret}}(+)=\boldsymbol{U}_{\alpha}^{\text{adv/ret}}(+)\boldsymbol{\Lambda}_{\alpha}^{\text{adv/ret}}(+)$ (20a)
$\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(-)^{-1}\boldsymbol{U}_{\alpha}^{\text{adv/ret}}(-)=\boldsymbol{U}_{\alpha}^{\text{adv/ret}}(-)\boldsymbol{\Lambda}_{\alpha}^{\text{adv/ret}}(-)^{-1}$ (20b)
@@
Appendix B Generalization to 2 or 3 dimensions
Although the examples given in this paper are only applied to 1D (monotatomic chain) or pseudo-1D systems (e.g. carbon nanotube), the generalization of our method to 2-dimensional (2D) or 3-dimensional (3D) systems merits a more detailed discussion. It should be noted that the original AGF method (Zhang et al., 2007b) has been applied to the investigation of phonon transport in 2D and 3D systems (Zhang et al., 2007a; Singh et al., 2011; Sadasivam et al., 2014) and the its extension, as described in Sec. A, can be applied in a straightforward manner. The key difference between a 1D system and its 2D or 3D counterpart is that in the latter, each phonon eigenmode of the lead, from Eq. (20b), is characterized by its transverse momentum ($k_{y},k_{z})$ in addition to its longitudinal momentum $k_{x}$, frequency $\omega$ and branch index $\nu$. Examples of the characterization of lead phonon eigenmodes by their transverse momenta can be found in Refs. (Ong, 2018b; Ong et al., 2020; Song and Chen, 2021; Ong, 2024).

### v2

@@ IV Summary and conclusions
In summary, we have shown rigorously how the mode-dependent scattering cross sections and rates in a lattice can be determined using the extended AGF method. (Ong, 2018a, b) This AGF-based approach allows us to treat scattering problems in which additional degrees of freedom from an external scatterer are present. Its utility is demonstrated with two examples: (1) the monoatomic harmonic chain with a single isotopic impurity, and (2) the (10,10) CNT with an encapsulated C60 molecule. In the first example, we derive the exact analytical expressions for the scattering cross section and rate. To first order in perturbation, the latter is identical to the result in Ref. (Allen and Nghiem, 2022). In the second example, we illustrate how the scattering cross section varies with the properties of the phonon mode (phonon branch and frequency), and show how phonon modes with radial atomic motion are more strongly scattered by the C60 molecule.
@@ Appendix A The extended Atomistic Green’s Function method
where $\boldsymbol{H}_{\text{C}}$, and $\boldsymbol{H}_{\text{CL}}$ ($\boldsymbol{H}_{\text{CR}}$) are respectively the mass-normalized force-constant submatrices corresponding to the device region, and the coupling between the device region and the semi-infinite left (right) lead. We can associate each layer in Fig. 1 with a block row in $\mathbf{H}$. The block row submatrices $\boldsymbol{H}_{\alpha}^{00}$ and $\boldsymbol{H}_{\alpha}^{01}$, where $\alpha=\text{L}$ and $\alpha=\text{R}$ for the left and right lead, respectively, describe the lead phonons. If we set the layers to be large enough so that only adjacent layers can couple, then $\boldsymbol{H}_{\alpha}^{00}$ denotes to the intralayer force-constant matrix for each layer while $\boldsymbol{H}_{\alpha}^{01}$ ($\boldsymbol{H}_{\alpha}^{10}$) denotes the interlayer harmonic coupling between each layer and the layer to its right (left) in the lead. Here, $\alpha$ is used as the dummy variable for distinguishing the leads, with $\alpha=\text{L}$ and $\alpha=\text{R}$ corresponding to the left and right leads, respectively. We also maintain the conceptual distinction between the left-lead and right-lead force-constant matrices of $\boldsymbol{H}_{\alpha}^{00}$ and $\boldsymbol{H}_{\alpha}^{01}$ even though $\boldsymbol{H}_{\text{L}}^{00}=\boldsymbol{H}_{\text{R}}^{00}$ and $\boldsymbol{H}_{\text{L}}^{01}=\boldsymbol{H}_{\text{R}}^{01}$ in our method for computing the scattering cross section.
@@
In principle, the system dynamics are determined by the infinitely large $\mathbf{H}$ in Eq. (15). However, for the effective dynamics at a fixed frequency $\omega$, the lattice dynamics problem becomes more tractable as we need only to project the dynamics onto a finite portion of the system, (Wang et al., 2008; Mingo, 2009) corresponding to layers $n=0$ to $N$ in Fig. 1, to determine phonon scattering by the device region. Hence, we can use the submatrices in Eq. (15) to construct the effective harmonic matrix for this subsystem, (Wang et al., 2008)
@@
where $\boldsymbol{H}_{\text{L}}^{\prime}=\boldsymbol{H}_{\text{L}}^{00}+\boldsymbol{H}_{\text{L}}^{10}\boldsymbol{g}_{\text{L},-}^{\text{ret}}\boldsymbol{H}_{\text{L}}^{01}$ and $\boldsymbol{H}_{\text{R}}^{\prime}=\boldsymbol{H}_{\text{R}}^{00}+\boldsymbol{H}_{\text{R}}^{01}\boldsymbol{g}_{\text{R},+}^{\text{ret}}\boldsymbol{H}_{\text{R}}^{10}$ correspond to the left and right edges, respectively while $\boldsymbol{H}_{\text{C}}^{\prime}=\boldsymbol{H}_{\text{C}}$ and $\boldsymbol{H}_{\text{CL/CR}}^{\prime}=\boldsymbol{H}_{\text{CL/CR}}=(\boldsymbol{H}_{\text{LC/RC}}^{\prime})^{\dagger}$. The retarded surface Green’s functions $\boldsymbol{g}_{\text{L},-}^{\text{ret}}(\omega)$ and $\boldsymbol{g}_{\text{R},+}^{\text{ret}}(\omega)$ are given by
$\boldsymbol{g}_{\alpha,-}^{\text{ret}}=[(\omega^{2}+i0^{+})\boldsymbol{I}_{\alpha}-\boldsymbol{H}_{\alpha}^{00}-\boldsymbol{H}_{\alpha}^{10}\boldsymbol{g}_{\alpha,-}^{\text{ret}}\boldsymbol{H}_{\alpha}^{01}]^{-1}\ ,$ (17a)
$\boldsymbol{g}_{\alpha,+}^{\text{ret}}=[(\omega^{2}+i0^{+})\boldsymbol{I}_{\alpha}-\boldsymbol{H}_{\alpha}^{00}-\boldsymbol{H}_{\alpha}^{01}\boldsymbol{g}_{\alpha,+}^{\text{ret}}\boldsymbol{H}_{\alpha}^{10}]^{-1}\ .$ (17b)
Physically, Eq. (17a) denotes the retarded surface Green’s function for a decoupled semi-infinite lattice extending infinitely to the left (denoted by the “-” in the subscript of $\boldsymbol{g}_{\alpha,-}^{\text{ret}}$) while Eq. (17b) denotes the corresponding surface Green’s function for a decoupled semi-infinite lattice extending infinitely to the right (denoted by the “+” in the subscript of $\boldsymbol{g}_{\alpha,+}^{\text{ret}}$). In addition, the advanced surface Green’s functions can be obtained from the Hermitian conjugates of Eq. (17b), i.e., $\boldsymbol{g}_{\alpha,-}^{\text{adv}}=(\boldsymbol{g}_{\alpha,-}^{\text{ret}}){}^{\dagger}$ and $\boldsymbol{g}_{\alpha,+}^{\text{adv}}=(\boldsymbol{g}_{\alpha,+}^{\text{ret}})^{\dagger}$.
@@
The next step is the computation of the advanced and retarded Bloch matrices (Ando, 1991; Khomyakov et al., 2005; Ong and Zhang, 2015) of the left and right leads, $\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(+)$ and $\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(-)$, which describe the bulk translational symmetry of the layers along the direction of the heat flux and can be computed directly from the formulas:
$\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(+)=\boldsymbol{g}_{\alpha,+}^{\text{adv/ret}}\boldsymbol{H}_{\alpha}^{10}\ ,$ (19a)
$\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(-)^{-1}=\boldsymbol{g}_{\alpha,-}^{\text{adv/ret}}\boldsymbol{H}_{\alpha}^{01}\ .$ (19b)
The bulk eigenmodes for the lead, which describe the mode-dependent atomic displacement in each layer at frequency $\omega$, can be determined directly from the Bloch matrices,
$\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(+)\boldsymbol{U}_{\alpha}^{\text{adv/ret}}(+)=\boldsymbol{U}_{\alpha}^{\text{adv/ret}}(+)\boldsymbol{\Lambda}_{\alpha}^{\text{adv/ret}}(+)\ ,$ (20a)
$\boldsymbol{F}_{\alpha}^{\text{adv/ret}}(-)^{-1}\boldsymbol{U}_{\alpha}^{\text{adv/ret}}(-)=\boldsymbol{U}_{\alpha}^{\text{adv/ret}}(-)\boldsymbol{\Lambda}_{\alpha}^{\text{adv/ret}}(-)^{-1}\ ,$ (20b)
@@
Appendix B Generalization to two or three dimensions
Although the examples given in this paper are only applied to 1D (monotatomic chain) or pseudo-1D systems (e.g. carbon nanotube), the generalization of our method to 2D or 3D systems merits a more detailed discussion. It should be noted that the original AGF method (Zhang et al., 2007b) has been applied to the investigation of phonon transport in 2D and 3D systems (Zhang et al., 2007a; Singh et al., 2011; Sadasivam et al., 2014) and its extension, as described in Appendix A, can be applied in a straightforward manner. The key difference between a 1D system and its 2D or 3D counterpart is that in the latter, each phonon eigenmode of the lead, from Eq. (20b), is characterized by its transverse momentum ($k_{y},k_{z})$ in addition to its longitudinal momentum $k_{x}$, frequency $\omega$ and branch index $\nu$. Examples of the characterization of lead phonon eigenmodes by their transverse momenta can be found in Refs. (Ong, 2018b; Ong et al., 2020; Song and Chen, 2021; Ong, 2024).